27y-8(dy/DX)³=0

1
Expert's answer
2022-06-10T05:31:50-0400
"27y-8(dy\/dx)^3=0"

"dy/dx=\sqrt[3]{27y/8}=3\sqrt[3]{y}/2"

"\frac{dy}{\sqrt[3]{y}}=\frac{3}{2}dx"

"\frac{3}{2}y^{2/3}=\frac{3}{2}(x+C)"

"y=(x+C)^{3\/2}"